OpenStudy (anonymous):

The population of a town in 2000 was 430. The population is increasing at a rate of 0.9% every year. What will be the projected population of the town in 2010? Round your answer to the nearest whole number

OpenStudy (campbell_st):

well the formula needed is similat to compound interest \[A = P(1 + \frac{r}{100})^t\] you have P = 430, r = 0.9 and t = 10 substitute and calculate